OpenStudy (anonymous):

What is the slope of the line that contains the points (-2, 8) and (-2, 1)?

hartnn (hartnn):

the slope of line is given by \(\large m=\dfrac{y_2-y_1}{x_2-x_1}\) here, x1,y1= -2,8 x2,y2=-2,1

OpenStudy (anonymous):

what?

hartnn (hartnn):

first find difference in y co-ordinates, what is y2-y1 = 1-8 = ..?

OpenStudy (anonymous):

i get it now thanks

hartnn (hartnn):

oh, welcome ^_^ what did u get the slope as ?